Problem Solving Using Systems of Equations
I purchased tickets to an air show for 5 adults and 2 children. The total cost was 162. The cost of a child's ticket was $3 less than the cost of an adult's ticket. Find the price of an adult's ticket and a child's ticket.

Answer by minerva(22)

cost of 1 adult ticket=xso 5 tickets=5x
cost of 1 child ticket =x-3 so 2 =2(x-3)
so,
2(x-3)+5x=162
2x-6+5x=162
7x=162+6
x=168/7
x=24
adult ticket =24
child ticket=24-3
=21